# Approximation by continuous functions and its applications

### v1

@@ Abstract
We prove that for every normal topological space $X$ and any function      $f:X\to\mathbb{R}$ there is a continuous function $g:X\to\mathbb{R}$ such that
$|f(x)-g(x)|\leq\frac{1}{2}\sup\limits_{p\in X}\inf\limits_{O(p)}\sup\limits_{a,b\in O(p)}|f(a)-f(b)|$
for all $x\in X$.
As an application of this result we prove the following statements to types of tightness in a space $Q_{p}(X,\mathbb{R})$ of all quasicontinuous real-valued functions with the topology $\tau_{p}$ of pointwise convergence: the countability of tightness (fan-tightness, strong fan-tightness) at a point $f$ of space $Q_{p}(X,\mathbb{R})$ implies the countability of tightness (fan-tightness, strong fan-tightness) of space $Q_{p}(X,Y)$ of all quasicontinuous functions from $X$ into any non-one-point metrizable space $Y$. This result is the answer to the open question in the class of metrizable spaces.
@@
approximation, quasicontinuous function, Lusin space, open Whyburn space, tightness, fan-tightness, strong fan-tightness, selection principle, continuous selection
@@ 1 Introduction
It is an open problem whether the space $Q_{p}(X,\mathbb{R})$ is topologically homogeneous (note that this space is not a linear space, since the sum of quasi-continuous functions is not necessarily quasi-continuous). In particular, it is unknown whether the points of $Q_{p}(X,\mathbb{R})$ can be distinguished using some topological property or cardinal invariant. In this regard, in (Osipov , Question 1) the second author raised a question on tightness in the space $Q_{p}(X,\mathbb{R})$. It was proved in Osipov that for every continuous $f:X\to\mathbb{R}$ we have $t(f,Q_{p}(X,\mathbb{R}))=\omega$ if and only if the space $X$ is $\mathcal{K}_{\Omega}$-Lindelof, and it was unknown whether the same is true for non-continuous functions.
@@
One of the main tools of our proof is Theorem 3.7 below on approximation of arbitrary functions by continuous functions (in the metric of uniform convergence). Although its formulation looks quite classical, the authors were unable to find this statement in the literature. Section 3 is devoted to the proof of this theorem.
In Sections 4 and 5, we prove some other technical lemmas needed for Section 6, where we finally investigate the properties of countable tightness, fan-tightness and strong fan-tightness in $Q_{p}(X,Y)$.
@@ 2 Preliminaries
(4) A space $X$ is said to be open Whyburn if for every open set $A\subset X$ and every $x\in\overline{A}\setminus A$ there is an open set $B\subseteq A$ such that $\overline{B}\setminus A=\{x\}$ Os1 .
Note that the class of open Whyburn spaces is quite wide; for example, it includes all first countable regular spaces Os1 and, therefore, all metrizable spaces.
@@
A space is $\mathcal{K}_{\Omega}$-Lindelöf if each element of $\mathcal{K}_{\Omega}$ has a countable subset in $\mathcal{K}_{\Omega}$ Sh2 .
@@
(Kun ) A Hausdorff space $X$ is called a Lusin space (in the sense of Kunen) if
@@
If $X$ is an uncountable Hausdorff space then $X$ is $\mathcal{O}^{s}$-Lindelöf (semi-Lindelöf) if and only if $X$ is a Lusin space (Corollary 2.5 in GJR ).
If $X$ is a Lusin space, $X$ is hereditarily Lindelöf (Lemma 1.2 in Kun ). Hence, if $X$ is a regular Lusin space then $X$ is perfect normal (3.8.A.(b) in Eng ).
If $X$ is a Lusin space, so is every uncountable subspace (Lemma 1.1 in Kun ).
Many topological properties are defined or characterized in terms of the following classical selection principles (see H1 ). Let $\mathcal{A}$ and $\mathcal{B}$ be sets consisting of families of subsets of an infinite set $X$. Then:
@@
For other notation and terminology almost without exceptions we follow the Engelking’s book Eng and the papers Os1 ; Osipov .
@@ 3 Approximation by continuous functions
Our goal in this section is to prove Theorem 3.7 and its corollary Theorem 3.9 below. The notion $\mathrm{dc}$ in its statement is defined as follows.
Definition 3.1.
Let $X$ be a topological space and let $(Y,d)$ be a metric space. For all functions $f:X\to Y$ we use the following notation:
  • 1.
    $\mathrm{dc}(f,E)=\sup\limits_{x,y\in E}d(f(x),f(y))$ for all $E\subseteq X$;
  • 2.
    $\mathrm{dc}(f,x)=\inf\limits_{O(x)}\mathrm{dc}(f,O(x))$;
  • 3.
    $\mathrm{dc}(f)=\sup\limits_{x\in X}\mathrm{dc}(f,x)$.
Our main tool in this section is Michael’s Selection Theorem. Before we formulate it, we must recall a few notions from HHM .
Definition 3.2.
Suppose that $X$ and $Y$ are topological spaces. A set-valued mapping $\Phi:X\to 2^{Y}$ is called lower semicontinuous, if for every $x_{0}\in X$, $y_{0}\in\Phi(x_{0})$ and a neighborhood $V$ of $y_{0}$ there is a neighborhood $U$ of $x_{0}$ such that for all $x\in U$ we have $\Phi(x)\cap V\neq\emptyset$.
Definition 3.3.
Let $X$ and $Y$ be topological spaces and let $\Phi:X\to 2^{Y}$ be a set-valued mapping. A function $f:X\to Y$ is called a selection of $\Phi$, if $f(x)\in\Phi(x)$ for all $x\in X$.
Theorem 3.4 (Modification of Michael’s Selection Theorem (Th. 2 in RS )).
The following properties of a $T_{1}$-space $X$ are equivalent:
(a) $X$ is normal;
(b) The real line $\mathbb{R}$ is an extensor for $X$;
(c) Every separable Banach space is extensor for $X$;
(d) Every lower semicontinuous set-valued mapping $\Phi:X\to\mathcal{C}(\mathbb{R})$ admits a continuous selection;
(e) If $Y$ is a separable Banach space, then every lower semicontinuous set-valued mapping $\Phi:X\to\mathcal{C}(Y)$ admits a continuous selection.
Here $\mathcal{C}(Y)$ denotes the family of all nonempty compact convex subsets of $Y$ or the whole space $Y$.
Now we only have to construct a suitable set-valued mapping.
Definition 3.5.
Suppose we are given a topological space $X$ and a function $f:X\to\mathbb{R}$. For all $x\in X$ we use the following notation:
  • 1.
    $f^{+}:X\to\mathbb{R}$ is the function
    $f^{+}(x)=\inf\limits_{O(x)}\sup\limits_{y\in O(x)}f(x);$
  • 2.
    $f^{-}:X\to\mathbb{R}$ is the function
    $f^{-}(x)=\sup\limits_{O(x)}\inf\limits_{y\in O(x)}f(x);$
  • 3.
    $\Phi_{f}(x)=[f^{+}(x)-\frac{1}{2}\mathrm{dc}(f),f^{-}(x)+\frac{1}{2}\mathrm{dc}(f)]$.
It is easy to see that $f^{+}(x)-f^{-}(x)=\mathrm{dc}(f,x)$ and $\Phi(x)\neq\emptyset$.
Lemma 3.6.
For any topological space $X$ and function $f:X\to\mathbb{R}$ the function $\Phi_{f}:X\to 2^{\mathbb{R}}$ is lower semicontinuous.
Proof.
Let us take any $x_{0}\in X$, any $y_{0}\in\Phi_{f}(x_{0})$ and any $\varepsilon>0$. There is a neighborhood $U$ of the point $x_{0}$ such that for every $x\in U$ we have
$f^{+}(x)<f^{+}(x_{0})+\varepsilon\text{ and }f^{-}(x)>f^{-}(x_{0})-\varepsilon.$
It follows that
$f^{+}(x)-\frac{1}{2}\mathrm{dc}(f)<f^{+}(x_{0})-\frac{1}{2}\mathrm{dc}(f)+\varepsilon\leq y_{0}+\varepsilon$
and
$f^{-}(x)+\frac{1}{2}\mathrm{dc}(f)<f^{-}(x_{0})+\frac{1}{2}\mathrm{dc}(f)-\varepsilon\leq y_{0}-\varepsilon,$
which means that $(y_{0}-\varepsilon,y_{0}+\varepsilon)\cap\Phi_{f}(x)\neq\emptyset$. ∎
@@
If $X$ is a normal space, then for every function $f:X\to\mathbb{R}$ there is a continuous function $g:X\to\mathbb{R}$ such that
$|f(x)-g(x)|\leq\frac{\mathrm{dc}(f)}{2}$
for all $x\in X$.
Proof.
We apply modification of Michael’s Selection Theorem to $\Phi_{f}$ and obtain some continuous selection $g$. It is easy to see that $|f(x)-g(x)|\leq\frac{1}{2}\mathrm{dc}(f)$ for all $x\in X$. ∎
Remark 3.8.
It is obvious that the upper bound $\frac{1}{2}\mathrm{dc}(f)$ is the best possible.
@@
Suppose $f(x)=\{f_{n}(x)\}_{n\in\mathbb{N}}$. For every $n\in\mathbb{N}$ choose some continuous function $g_{n}:X\to\mathbb{R}$ such that $|f_{n}(x)-g_{n}(x)|\leq\frac{\mathrm{dc}(f_{n})}{2}\leq\frac{\mathrm{dc}(f)}{2}$ for all $x\in X$. Define $g(x)=\{g_{n}(x)\}_{n\in\mathbb{N}}$ for all $x\in X$. It is easy to see that $g:X\to\mathbb{H}$ is continuous and $d_{\mathbb{H}}(f(x),g(x))\leq\frac{\mathrm{dc}(f)}{2}$. ∎
@@ 4 Domain approximation
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. We say that a countable family $\mathcal{E}$ of closed subsets of $X$ is a $\mathrm{DA}$-family (or domain approximation family) for the function $f$ if for every finite $K\subseteq X$ and any $\varepsilon>0$ there is a set $F\in\mathcal{E}$ such that $K\subseteq F$ and $\mathrm{dc}(f|_{F})<\varepsilon$.
@@
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. For all $\varepsilon>0$ we denote $D_{\varepsilon}(f)=\{x\in X:\mathrm{dc}(f,x)\geq\varepsilon\}$. We also denote $D(f)=\bigcup\limits_{\varepsilon>0}D_{\varepsilon}(f)$, which is the set of all discontinueties of $f$.
@@
If $X$ is a metrizable Luzin space, then for every metric space $Y$ and any $f\in Q(X,Y)$ there is a $\mathrm{DA}$-family for $f$.
@@ 5 Flattening lemma
Recall that the $\omega$-th power of the space $H_{S}$ is an universal space for all metrizable spaces of weight $\leq|S|$ (see Eng ). We fix on the space $H_{S}^{\omega}$ the metric
@@
We can suppose that $X\subseteq H_{S}^{\omega}$ for some uncountable set $S$ (Eng, , Theorem 4.4.9). Since $A$ is separable, we can choose a countable subset $C\subseteq S$ such that $A\subseteq H_{C}^{\omega}$. Choose any $\gamma\in S\setminus C$ and denote $D=C\cup\{\gamma\}$. Now for all $x\in X$ define $h(x)=\{h_{D}^{\gamma}(x(n))\}_{n\in\omega}$, so $h:X\to H_{D}^{\omega}$. It follows from the previous propostion that $\rho_{S}^{\omega}(h(x),h(y))=\rho_{S}^{\omega}(x,y)$ for all $x\in A$ and $y\in X$. Finally, we take any homeomorphic embedding $g:H_{D}\to\mathbb{H}$ and define $\varphi(x)=g(h(x))$ for all $x\in X$. The function $\varphi$ is as required. ∎
@@ 6.1 Tightness
For every $h\in P$ and $n\in\mathbb{N}$ denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in Lev ).
@@
(2) Let $IntF\neq\emptyset$. By Theorem 3.6 ($(2)\Rightarrow(5)(a)$) in Osipov (Any open subset of $\mathcal{K}_{\Omega}$-Lindelöf space is $\mathcal{K}_{\Omega}$-Lindelöf), $IntF$ is $\mathcal{K}_{\Omega}$-Lindelöf and, by Theorem 3.6 ($(2)\Rightarrow(5)(b)$) (The union of $\mathcal{K}_{\Omega}$-Lindelöf space with countable space is $\mathcal{K}_{\Omega}$-Lindelöf), $F=IntF\cup(F\setminus IntF)$ is $\mathcal{K}_{\Omega}$-Lindelöf.
@@
$(4)\Rightarrow(1)$. By Lemma 3.1 in Osipov , $X$ is a Lusin space.
@@ 6.2 Fan-tightness
For every $n\in\mathbb{N}$ and $h\in P_{n}$ and denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in Lev ).
@@
$(3)\Rightarrow(1)$. It is clear that any space satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$ is $\mathcal{K}_{\Omega}$-Lindelöf. Then, by Lemma 3.1 in Osipov , $X$ is a Lusin space.
@@ 6.3 Strong fan-tightness
Combining the results in Os1 ; Osipov and this paper, we get the following diagram in the class of metrizable spaces $X$ and $Y$.
@@
In (Osipov , Example 6.1) it was shown that, under Jensen’s axiom, there exists a metrizable Lusin space $X$ such that $vet_{1}(Q_{p}(X,\mathbb{R}))=\omega$. An interesting open question (Questions 2 and 3 in Osipov ) remains: Is there a $T_{2}$-space $X$ such that $Q_{p}(X,\mathbb{R})$ has countable tightness but none $Q_{p}(X,\mathbb{R})$ has countable (strong) fan-tightness?

### v2

# Tightness type properties of spaces of quasicontinuous functions

@@ Abstract
Using approximation by continuous functions we prove the following statements to types of tightness in a space $Q_{p}(X,\mathbb{R})$ of all quasicontinuous real-valued functions with the topology $\tau_{p}$ of pointwise convergence: the countability of tightness (fan-tightness, strong fan-tightness) at a point $f$ of space $Q_{p}(X,\mathbb{R})$ implies the countability of tightness (fan-tightness, strong fan-tightness) of space $Q_{p}(X,Y)$ of all quasicontinuous functions from $X$ into any non-one-point metrizable space $Y$. This result is the answer to the open question in the class of metrizable spaces.
@@
quasicontinuous function , Lusin space , open Whyburn space , tightness , fan-tightness , strong fan-tightness , selection principle , continuous selection
@@ 1 Introduction
It is an open problem whether the space $Q_{p}(X,\mathbb{R})$ is topologically homogeneous (note that this space is not a linear space, since the sum of quasi-continuous functions is not necessarily quasi-continuous). In particular, it is unknown whether the points of $Q_{p}(X,\mathbb{R})$ can be distinguished using some topological property or cardinal invariant. In this regard, in ([5], Question 1) the second author raised a question on tightness in the space $Q_{p}(X,\mathbb{R})$. It was proved in [5] that for every continuous $f:X\to\mathbb{R}$ we have $t(f,Q_{p}(X,\mathbb{R}))=\omega$ if and only if the space $X$ is $\mathcal{K}_{\Omega}$-Lindelof, and it was unknown whether the same is true for non-continuous functions.
@@
One of the main tools of our proof is Theorem 3.1 below on approximation of arbitrary functions by continuous functions (in the metric of uniform convergence).
@@ 2 Preliminaries
(4) A space $X$ is said to be open Whyburn if for every open set $A\subset X$ and every $x\in\overline{A}\setminus A$ there is an open set $B\subseteq A$ such that $\overline{B}\setminus A=\{x\}$ [4].
Note that the class of open Whyburn spaces is quite wide; for example, it includes all first countable regular spaces [4] and, therefore, all metrizable spaces.
@@
A space is $\mathcal{K}_{\Omega}$-Lindelöf if each element of $\mathcal{K}_{\Omega}$ has a countable subset in $\mathcal{K}_{\Omega}$ [9].
@@
([3]) A Hausdorff space $X$ is called a Lusin space if
@@
If $X$ is an uncountable Hausdorff space then $X$ is $\mathcal{O}^{s}$-Lindelöf (semi-Lindelöf) if and only if $X$ is a Lusin space (Corollary 2.5 in [10]).
If $X$ is a Lusin space, $X$ is hereditarily Lindelöf (Lemma 1.2 in [3]). Hence, if $X$ is a regular Lusin space then $X$ is perfect normal (3.8.A.(b) in [2]).
If $X$ is a Lusin space, so is every uncountable subspace (Lemma 1.1 in [3]).
Many topological properties are defined or characterized in terms of the following classical selection principles (see [8]). Let $\mathcal{A}$ and $\mathcal{B}$ be sets consisting of families of subsets of an infinite set $X$. Then:
@@
For other notation and terminology almost without exceptions we follow the Engelking’s book [2] and the papers [4, 5].
@@ 3 Approximation by continuous functions
Let $X$ be a normal space. For any $f\in\mathbb{R}^{X}$ there exists a continuous function $g:X\rightarrow\mathbb{R}$ such that $|f(x)-g(x)|\leq\frac{osc(f)}{2}$, where $osc(f)$ is the oscillation of $f$.
This result can be found, for instance, in [1] (Proposition 1.18) (see also [2], Problem 1.7.15 (b)).
@@
Suppose $f(x)=\{f_{n}(x)\}_{n\in\mathbb{N}}$. For every $n\in\mathbb{N}$ choose some continuous function $g_{n}:X\to\mathbb{R}$ such that $|f_{n}(x)-g_{n}(x)|\leq\frac{osc(f_{n})}{2}\leq\frac{osc(f)}{2}$ for all $x\in X$. Define $g(x)=\{g_{n}(x)\}_{n\in\mathbb{N}}$ for all $x\in X$. It is easy to see that $g:X\to\mathbb{H}$ is continuous and $d_{\mathbb{H}}(f(x),g(x))\leq\frac{osc(f)}{2}$. ∎
@@ 4 Domain approximation
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. We say that a countable family $\mathcal{E}$ of closed subsets of $X$ is a $\mathrm{DA}$-family (or domain approximation family) for the function $f$ if for every finite $K\subseteq X$ and any $\varepsilon>0$ there is a set $F\in\mathcal{E}$ such that $K\subseteq F$ and $osc(f|_{F})<\varepsilon$.
@@
Suppose we are given a topological space $X$, a metric space $Y$ and a function $f:X\to Y$. For all $\varepsilon>0$ we denote $D_{\varepsilon}(f)=\{x\in X:osc(f,x)\geq\varepsilon\}$. We also denote $D(f)=\bigcup\limits_{\varepsilon>0}D_{\varepsilon}(f)$, which is the set of all discontinueties of $f$.
@@
If $X$ is a metrizable Lusin space, then for every metric space $Y$ and any $f\in Q(X,Y)$ there is a $\mathrm{DA}$-family for $f$.
@@ 5 Flattening lemma
Recall that the $\omega$-th power of the space $H_{S}$ is an universal space for all metrizable spaces of weight $\leq|S|$ (see [2]). We fix on the space $H_{S}^{\omega}$ the metric
@@
We can suppose that $X\subseteq H_{S}^{\omega}$ for some uncountable set $S$ [2, Theorem 4.4.9]. Since $A$ is separable, we can choose a countable subset $C\subseteq S$ such that $A\subseteq H_{C}^{\omega}$. Choose any $\gamma\in S\setminus C$ and denote $D=C\cup\{\gamma\}$. Now for all $x\in X$ define $h(x)=\{h_{D}^{\gamma}(x(n))\}_{n\in\omega}$, so $h:X\to H_{D}^{\omega}$. It follows from the previous propostion that $\rho_{S}^{\omega}(h(x),h(y))=\rho_{S}^{\omega}(x,y)$ for all $x\in A$ and $y\in X$. Finally, we take any homeomorphic embedding $g:H_{D}\to\mathbb{H}$ and define $\varphi(x)=g(h(x))$ for all $x\in X$. The function $\varphi$ is as required. ∎
@@ 6.1 Tightness
For every $h\in P$ and $n\in\mathbb{N}$ denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in [6]).
@@
(2) Let $IntF\neq\emptyset$. By Theorem 3.6 ($(2)\Rightarrow(5)(a)$) in [5] (Any open subset of $\mathcal{K}_{\Omega}$-Lindelöf space is $\mathcal{K}_{\Omega}$-Lindelöf), $IntF$ is $\mathcal{K}_{\Omega}$-Lindelöf and, by Theorem 3.6 ($(2)\Rightarrow(5)(b)$) (The union of $\mathcal{K}_{\Omega}$-Lindelöf space with countable space is $\mathcal{K}_{\Omega}$-Lindelöf), $F=IntF\cup(F\setminus IntF)$ is $\mathcal{K}_{\Omega}$-Lindelöf.
@@
$(4)\Rightarrow(1)$. By Lemma 3.1 in [5], $X$ is a Lusin space.
@@ 6.2 Fan-tightness
For every $n\in\mathbb{N}$ and $h\in P_{n}$ and denote by $V_{h,n}=\{x\in X:d(f(x),h(x))<\varepsilon+\frac{1}{n}\}$. Note that $V_{h,n}$ is semi-open because $h\in Q(X,Y)$ and $V_{h,n}=\bigcup\{h^{-1}(\{y\in Y:d(f(x),y)<\varepsilon+\frac{1}{n}\}):x\in X\}$ (see Theorem 2 in [6]).
@@
$(3)\Rightarrow(1)$. It is clear that any space satisfies $S_{\mathrm{fin}}(\mathcal{K}_{\Omega},\mathcal{K}_{\Omega})$ is $\mathcal{K}_{\Omega}$-Lindelöf. Then, by Lemma 3.1 in [5], $X$ is a Lusin space.
@@ 6.3 Strong fan-tightness
Combining the results in [4, 5] and this paper, we get the following diagram in the class of metrizable spaces $X$ and $Y$.
@@
In ([5], Example 6.1) it was shown that, under Jensen’s axiom, there exists a metrizable Lusin space $X$ such that $vet_{1}(Q_{p}(X,\mathbb{R}))=\omega$. An interesting open question (Questions 2 and 3 in [5]) remains:
Is there a $T_{2}$-space $X$ such that $Q_{p}(X,\mathbb{R})$ has countable tightness but none $Q_{p}(X,\mathbb{R})$ has countable (strong) fan-tightness?